solve absolute value of x - 4 = 5 - 2x
Answer by stanbon(48546)

solve |x - 4| = 5 - 2x
x-4 = 5-2x or 4-x = 5-2x
3x = 9 or x = 1
x = 3 or x = 1
===================
Cheers,
Stan H.